OpenStudy (anonymous):

I need help with these steps please. ( x^3+2x^2-3x+2)÷(x+1)

OpenStudy (anonymous):

do you know synthetic division?

OpenStudy (gw2011):

The answer is: x^2+x-4 Remainder 6

OpenStudy (anonymous):

I need also the step please. ( x^3+2x^2-3x+2)÷(x+1)

OpenStudy (gw2011):

x divided into x^3 is x^2 x^2 times x+1 = x^3+x^2 x^3+x^2 subtracted from x^3+2x^2-3x+2 = x^2-3x+2 x divided into x^2 is x x times x+1 = x^2+x x^2+x subtracted from x^2-3x+2 = -4x+2 x divided into-4x is -4 -4 times x+1 = -4x-4 -4x-4 subtracted from -4x+2 = 6 This is the remainder